40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Fri Oct 4 12:33:27 2013, Masayuki, Update, SUS, Replaced the laser for Optical Lever of ETMY 
|
Fri Oct 4 14:36:44 2013, rana, Update, SUS, Replaced the laser for Optical Lever of ETMY
|
Sun Oct 6 18:37:43 2013, rana, Update, SUS, Replaced the laser for Optical Lever of ETMY
|
Mon Oct 7 10:51:18 2013, Steve, Update, SUS, Replaced the laser for Optical Lever of ETMY 
|
Tue Oct 8 08:58:16 2013, Steve, Update, SUS, Oplev power spectrums  
|
Thu Oct 10 11:30:15 2013, Steve, Update, SUS, Oplev power spectrums: PIT-YAW
|
Tue Oct 15 15:25:13 2013, Steve, Update, SUS, Oplev power spectrums bandwidths 
|
Wed Oct 16 10:32:40 2013, Steve, Update, SUS, Flowbench effect on oplev error signals
|
Thu Jul 10 15:57:00 2014, Steve, Update, SUS, flow bench is turned off
|
Thu Oct 24 17:18:44 2013, Steve, Update, SUS, ETMY oplev error signals are still bad
|
|
Message ID: 9203
Entry time: Fri Oct 4 14:36:44 2013
In reply to: 9202
|
Author: |
rana |
Type: |
Update |
Category: |
SUS |
Subject: |
Replaced the laser for Optical Lever of ETMY |
|
|
That's good, but please no more Oplev work. We want to do all of it at once and to make no more changes until we have all the parts (e.g. dumps and correct lenses) in hand and then talk over what the new design will be. I don't want to tune the beam size and loop shape every week. |